OpenStudy (anonymous):

How do you solve 2/9 + 19/81+ 12/27

OpenStudy (muscrat123):

i think u get a common denominator

OpenStudy (wolf1728):

Make all the denominators of each fraction equal 81.

OpenStudy (wolf1728):

If you get the fraction 2/9, and changed it to ?/81, what would the numerator equal?

OpenStudy (anonymous):

take LCM of 9, 81, 27, \[\frac{ 2 }{ 9 }+\frac{ 19 }{ 81 }+\frac{ 12 }{ 27 } \\= \frac{ 2 }{ 9 }\times \frac{ 9 }{ 9 }+\frac{ 19 }{ 81 }\times \frac{ 1 }{ 1 }+\frac{ 12 }{ 27}\times \frac{ 3 }{ 3 }\\=\frac{ 18 }{ 81 }+\frac{ 19 }{ 81} + \frac{ 36 }{ 81} \\=\frac{ 18+19+36 }{ 81 }\\=\frac{ 73 }{ 81 }Ans\]